# To explain the muon excess phenomenon in cosmic rays based on the gluon condensation model

### v1

@@ 3.2 Production of pions and kaons in high-energy collisions
$\sigma_{gg}(s)=\frac{\pi\alpha_{s}^{2}}{3s}$ $\left[(1+\eta+\frac{1}{16}\eta^{2})\ln\left(\frac{1+\sqrt{1-\eta}}{1-\sqrt{1-\eta}}\right)\right.$ (11)
$\left.-\left(\frac{7}{4}+\frac{31}{16}\eta\right)\sqrt{1-\eta}\right],$
in the two equations, $s=\left(q+\bar{q}\right)^{2}$ or $s=\left(g+g\right)^{2}$ is the center-of-mass energy of the corresponding process. $\eta={4m_{s}^{2}}/s$, where $m_{s}$ is the mass of the strange quark and its antiquark. The reaction will not appear if the energy $\sqrt{s}$ is less than the threshold energy $2m_{s}$. In the QGP state, the momentum distribution functions of quarks and gluons follow the Fermi-Dirac and Bose-Einstein distributions, respectively.
@@
×$\times\left\{N_{c}^{2}N_{s}^{2}N_{f}\int\frac{\mathrm{d}^{3}k_{1}\mathrm{d}^{3}k_{2}}{(2\pi)^{6}|k_{1}||k_{2}|}f_{q}(k_{1})f_{\bar{q}}(k_{2})\sigma_{q\bar{q}}(s)\right.$
$\left.+N_{g}^{2}N_{\epsilon}^{2}\int\frac{\mathrm{d}^{3}k_{1}\mathrm{d}^{3}k_{2}}{(2\pi)^{6}|k_{1}||k_{2}|}f_{g}(k_{1})f_{g}(k_{2})\sigma_{gg}(s)\right\}.$
where $N_{c}=3$ is the color number of quarks, $N_{g}=8$ is the color number of gluons, $N_{\epsilon}=2$ is the number of polarizations, $N_{f}=2$ is the flavor degeneracy in QGP. For the process $q+\bar{q}\rightarrow s+\bar{s}$, we have $N_{q\bar{q}}=N_{c}^{2}N_{s}^{2}N_{f}=72$; for the process $g+g\rightarrow s+\bar{s}$, we have $N_{gg}=N_{g}^{2}N_{\epsilon}^{2}=256$. It can be seen that the total production rate of strangeness density is mainly dominated by gluon collisions. Many experimental collaborations, such as CERN’s NA34 Van Hecke (1991); Mazzoni et al. (1992) and BNL’s E802 Nagamiya (1992); Abbott et al. (1991); Zajc (1991), have conducted research on strangeness in high-energy heavy-ion collisions. In condition of the incident particle energy at the scale of GeV, $n_{K^{+}}/n_{\pi^{+}}$ can reach approximately 0.24 Wong (1994). The production of pions and kaons in condition of quark-gluon plasma reaching thermal and chemical equilibrium at $T=200$ MeV, $n_{K^{+}}/n_{\pi^{+}}\approx 0.38$. Comparatively, with the increasing of the energy scale of the incident particles, the strangeness increase. While, it’s still not rich enough to solve the problem of muon excess even in conditions of QGP.
@@
In GC conditions, the distribution of gluons tends to be a delta function like, assuming
@@
where $k_{c}=\frac{3}{2}T$ and $C_{m}<1.0$, we take $C_{m}=0.4$ based on the QCD calculations and global fitting of gluon distributions Gao et al. (2018). In fact, the specific value of $C_{m}$ will not affect our final conclusion. Substituting Eq.15 into Eq.13, we obtain,
@@
It’s evident that the value of $A_{\mathrm{GC}}$ is significantly greater than $A_{\mathrm{QGP}}$ in Figure 3. $A_{\mathrm{GC}}/A_{\mathrm{OGP}}\approx 8.9$. The production rates of strange quarks in both conditions are positively correlated with temperature. Although the trends of the two curves appear similar in the plot, it is noticeable that the value of $A_{\mathrm{GC}}$ has been larger than $A_{\mathrm{QGP}}$ at $T=300$ MeV. This indicates that the production rate of strange quarks in GC condition is much richer than that in QGP condition. So, a significant enhancement in the number of kaons in the hadronization products can be expected in the gluon condensation state, which could potentially explain the muon excess phenomenon.
@@ 3.3 Productions of Pions and Kaons in the GC Model
In this section, we will introduce the production of pions and kaons in the gluon condensation model. In this model, the distribution of gluons at the critical momentum $k_{c}$ is shown in Eq.15. When the energy in the center of momentum system in collisions reaches the energy threshold, gluon condensation will occur. In our model, the first stage of air shower cascade, which initiated by ultrahigh-energy cosmic rays (taking $E_{0}=10^{19}$ eV as an example), reach the energy threshold for the occurrence of gluon condensation. For simplicity, we assume that only pions and kaons are the new productions. In condition of GC state, abundant strange quarks( $g+g\rightarrow s+\bar{s}$) are generated from the interactions between the huge number of gluons. In the subsequent hadronization processes, the number of kaons increases significantly. Kaons include $K^{+}$, $K^{-}$, $K^{0}_{L}$, and $K^{0}_{S}$. At the EeV energy scale, the isospin symmetry breaking in kaons can be ignored Adhikary et al. (2023b), so the four types of kaons in this paper are assumed to be produced equally. Except for kaons, large number of pions also are produced simultaneously.
@@
In this study, the theoretical framework outlined in Eq.4 to Eq.8 serves as the basis to simulate the secondary particle products resulting from showers in proton-proton collisions in various conditions. We consider three conditions seperatly, that is , the leading protons, carrying away proportions of the primary energy with $50\%$, $60\%$ and $70\%$, which means that the total energy carried by pions and kaons are $50\%$, $40\%$, and $30\%$ correspondingly. These simulations are compared with the traditional result derived from the SIBYLL model Fletcher et al. (1994); Engel et al. (2017). In Figure 4, the simulated results in the GC condition for pions and kaons are shown. They illustrate the relative increases of $N_{\mu}$ and $X_{\mathrm{max}}$ compared to the traditional model, along with the $n_{K}/n_{\pi}$. The relative increases are quantified as $(N_{\mu}-N_{\mu}^{\mathrm{SIBYLL}})/N_{\mu}^{\mathrm{SIBYLL}}$ and $(X_{\mathrm{max}}-X_{\mathrm{max}}^{\mathrm{SIBYLL}})/X_{\mathrm{max}}^{\mathrm{SIBYLL}}$. There are some typical characters of the results:
@@
(3) Since the SIBYLL model does not indicate obvious deviation between $X_{\mathrm{max}}$ and the experimental data, we can conclude that the pions and kaons carring $50\%$ of the total energy (Figure 4(b)) is not a suitable choice. In Figure 4(d) and (f), the relative deviation is within the range of $0.07$ and $0.04$. Considering the corresponding experimental error of $X_{\mathrm{max}}$, it is acceptable to take the values $40\%$ or $30\%$ of the primary energy carried by mesons as in Figure 4(d) and (f). A better balance can be achieved between solving the problem of muon excess and maintaining the consistency with the depth of shower maximum.
@@ 4 Discussion and conclusion
In high-energy collision experiments, the existing researches suggest that QGP may occur, but considering only the conventional QGP effect cannot solve the muon excess problem in air shower. We consider that GC may occur in extremely high energy collisions. In the GC state, the production of strange quarks will be greatly enhanced, so that the $n_{K}/n_{\pi}$ can far exceed the $n_{K}/n_{\pi}\approx 0.24$ in ordinary situations and $n_{K}/n_{\pi}\approx 0.38$ in QGP, reaching $n_{K}/n_{\pi}>1.0$. In figure.4 and figure.5, we simulated $(N_{\mu}-N_{\mu}^{\mathrm{SIBYLL}})/N_{\mu}^{\mathrm{SIBYLL}}$ and $(X_{\mathrm{max}}-X_{\mathrm{max}}^{\mathrm{SIBYLL}})/X_{\mathrm{max}}^{\mathrm{SIBYLL}}$ in the $1.00.1<n_{K}/n_{\pi}<$ 1.0 region under GC conditions and obtained reasonable results.
According to the researches in work Anchordoqui et al. (2017), when the energy of incident particle is about $10^{19}$ eV, the measured number of muons exceeds the theoretical prediction by approximately $40\%$. In this study, considering only the gluon condensation effect in the proton-proton collisions, the results in figure.4 indicate a significant enhancement in $N_{\mu}$ in air showers. For $n_{K}/n_{\pi}=0.4$ , the muon number in the GC model reaches the expected enhancement $40\%$ when the mesons taken about $30\%\sim 40\%$ of the incident energy. However, as pointed out by Albrecht et al. (2022), the muon excess phenomenon becomes more pronounced with the increasing primary energy. In GC model, considering that $n_{K}/n_{\pi}$ will increase as the the energy of incident cosmic rays increase, it will be naturally obtained.

### v2

@@ 3.2 Production of pions and kaons in high-energy collisions
$\sigma_{gg}(s)=\frac{\pi\alpha_{s}^{2}}{3s}\left[(1+\eta+\frac{1}{16}\eta^{2})\ln\left(\frac{1+\sqrt{1-\eta}}{1-\sqrt{1-\eta}}\right)-\left(\frac{7}{4}+\frac{31}{16}\eta\right)\sqrt{1-\eta}\right],$ (11)
in the two equations, $s=\left(q+\bar{q}\right)^{2}$ or $s=\left(g+g\right)^{2}$ is the square of center-of-mass energy of the corresponding process. $\eta={4m_{s}^{2}}/s$, where $m_{s}$ is the mass of the strange quark and its antiquark. The reaction will not appear if the energy $\sqrt{s}$ is less than the threshold energy $2m_{s}$. In the QGP state, the momentum distribution functions of quarks and gluons follow the Fermi-Dirac and Bose-Einstein distributions, respectively.
@@
×$\times\left\{N_{c}^{2}N_{s}^{2}N_{f}\int\frac{\mathrm{d}^{3}k_{1}\mathrm{d}^{3}k_{2}}{(2\pi)^{6}|k_{1}||k_{2}|}f_{q}(k_{1})f_{\bar{q}}(k_{2})\sigma_{q\bar{q}}(s)+N_{g}^{2}N_{\epsilon}^{2}\int\frac{\mathrm{d}^{3}k_{1}\mathrm{d}^{3}k_{2}}{(2\pi)^{6}|k_{1}||k_{2}|}f_{g}(k_{1})f_{g}(k_{2})\sigma_{gg}(s)\right\}.$
where $N_{c}=3$ is the color number of quarks, $N_{g}=8$ is the color number of gluons, $N_{\epsilon}=2$ is the number of polarizations, $N_{f}=2$ is the flavor degeneracy in QGP. For the process $q+\bar{q}\rightarrow s+\bar{s}$, we have $N_{q\bar{q}}=N_{c}^{2}N_{s}^{2}N_{f}=72$; for the process $g+g\rightarrow s+\bar{s}$, we have $N_{gg}=N_{g}^{2}N_{\epsilon}^{2}=256$. It can be seen that the total production rate of strangeness density is mainly dominated by gluon collisions. Many experimental collaborations, such as CERN’s NA34 Van Hecke (1991); Mazzoni et al. (1992) and BNL’s E802 Nagamiya (1992); Abbott et al. (1991); Zajc et al. (1992), have conducted research on strangeness in high-energy heavy-ion collisions. In condition of the incident particle energy at the scale of GeV, $n_{K^{+}}/n_{\pi^{+}}$ can reach approximately 0.24 Wong (1994). In condition of thermal and chemical equilibrium at $T=200$ MeV, the production of pions and kaons in quark-gluon plasma state will be $n_{K^{+}}/n_{\pi^{+}}\approx 0.38$. Although the strangeness increases, it’s still not rich enough to solve the problem of muon excess Anchordoqui et al. (2017).
@@
In GC conditions, the distribution of gluons tends to be a delta function like,
@@
where $k_{c}=\frac{3}{2}T$ and $C_{m}\textless 1.0$, we take $C_{m}=0.4$ (a lower limit) based on the QCD calculations and global fitting of gluon distributions Gao et al. (2018). Substituting Eq.15 into Eq.13, we obtain,
@@
It’s evident that the value of $A_{\mathrm{GC}}$ is significantly greater than $A_{\mathrm{QGP}}$ in Figure 3 with $A_{\mathrm{GC}}/A_{\mathrm{OGP}}\approx 8.9$. The production rates of strange quarks in both conditions are positively correlated with temperature. The trends of the two curves appear similar in the plot, and the value of $A_{\mathrm{GC}}$ shows much larger than $A_{\mathrm{QGP}}$ within the range of temperature $T$. This indicates that the production rate of strange quarks in GC condition is much richer than that in QGP condition. So, a significant enhancement in the number of kaons in the hadronization products can be expected in the gluon condensation state.
@@ 3.3 Productions of Pions and Kaons in the GC Model
In this section, we will introduce the production of pions and kaons in the gluon condensation model. In this model, the distribution of gluons at the critical momentum $k_{c}$ is shown in Eq.15. When the energy of colliding particles in the center of momentum system reaches the energy threshold, gluon condensation will occur. In our model, the first stage of air shower cascade, initiated by ultrahigh-energy cosmic rays (taking $E_{0}=10^{19}$ eV as an example), reach the energy threshold for the occurrence of gluon condensation. For simplicity, we assume that only pions and kaons are the new productions. In condition of GC state, abundant strange quarks are generated from the interactions between the huge number of gluons ($g+g\rightarrow s+\bar{s}$). In the subsequent hadronization processes, the number of kaons increases significantly. Kaons include $K^{+}$, $K^{-}$, $K^{0}_{L}$, and $K^{0}_{S}$. At the EeV energy scale, the isospin symmetry breaking in kaons can be ignored Adhikary et al. (2023b), so the four types of kaons in this paper are assumed to be produced equally.
@@
In this study, the theoretical framework outlined in Eq.4 to Eq.8 serves as the basis to simulate the secondary particle products resulting from showers in proton-proton collisions in various conditions. We consider three conditions seperatly, that is , the leading protons, carrying away proportions of the primary energy with $50\%$, $60\%$ and $70\%$, which means that the total energy carried by pions and kaons are $50\%$, $40\%$, and $30\%$ correspondingly. These simulations are compared with the traditional result derived from the SIBYLL model Fletcher et al. (1994); Engel et al. (2017). In Figure 4, the simulated results in the GC condition for pions and kaons are shown. They illustrate the relative increases of $N_{\mu}$ and $X_{\mathrm{max}}$ compared to the traditional model, along with the $n_{K}/n_{\pi}$. The relative increases are quantified as $(N_{\mu}-N_{\mu}^{\mathrm{SIBYLL}})/N_{\mu}^{\mathrm{SIBYLL}}$ and $(X_{\mathrm{max}}-X_{\mathrm{max}}^{\mathrm{SIBYLL}})/X_{\mathrm{max}}^{\mathrm{SIBYLL}}$. There are some key characteristics of the results:
@@
(3) Since the SIBYLL model does not indicate obvious deviation between simulated $X_{\mathrm{max}}$ and the experimental data, we can conclude that the pions and kaons carring $50\%$ of the total energy (Figure 4(b)) is not a suitable choice. In Figure 4(d) and (f), the relative deviation is within the range of $0.07$ and $0.04$. Considering the corresponding experimental error of $X_{\mathrm{max}}$, it is acceptable to take the values $40\%$ or $30\%$ of the primary energy carried by mesons as in Figure 4(d) and (f). A better balance can be achieved between solving the problem of muon excess and maintaining the consistency with the depth of shower maximum.
@@ 4 Discussion and conclusion
In high-energy collision experiments, the existing research suggest that QGP may occur, but considering only the conventional QGP effect cannot solve the muon excess problem in air shower. We consider that GC may occur in extremely high energy collisions. In the GC state, the production of strange quarks will be greatly enhanced, so that the $n_{K}/n_{\pi}$ can far exceed the $n_{K}/n_{\pi}\approx 0.24$ in ordinary situations and $n_{K}/n_{\pi}\approx 0.38$ in QGP, reaching $n_{K}/n_{\pi}>1.0$. In figure.4 and figure.5, we simulated $(N_{\mu}-N_{\mu}^{\mathrm{SIBYLL}})/N_{\mu}^{\mathrm{SIBYLL}}$ and $(X_{\mathrm{max}}-X_{\mathrm{max}}^{\mathrm{SIBYLL}})/X_{\mathrm{max}}^{\mathrm{SIBYLL}}$ in the $1.00.1<n_{K}/n_{\pi}<$ 1.0 region under GC conditions and obtained reasonable results.
According to Anchordoqui et al. (2017), when the energy of incident particle is about $10^{19}$ eV, the measured number of muons exceeds the theoretical prediction by approximately $40\%$. In this study, considering only the gluon condensation effect in the proton-proton collisions, the results in figure.4 indicate a significant enhancement in $N_{\mu}$ in air showers. The muon number increases with the value of $n_{K}/n_{\pi}$ increasing. For $n_{K}/n_{\pi}=0.6$, the production of muon in the GC model reaches the expected enhancement of $40\%$ when the mesons taken about $30\%\sim 40\%$ of the incident energy. When $n_{K}/n_{\pi}>0.6$, the increase becomes relatively slow. Considering that the results should not conflict with the measured $X_{\mathrm{max}}$ simultaneously, the number of muons in our model can be increased by up to $50\%$ compared to the traditional models in condition of $n_{K}/n_{\pi}<1.0$. As pointed out by Albrecht et al. (2022), the muon excess phenomenon becomes more pronounced with the increasing primary energy. In GC model, considering that $n_{K}/n_{\pi}$ will increase as the the energy of incident cosmic rays increase, it will be naturally obtained.
@@
Acknowledgements
This work is supported by the National Natural Science Foundation of China (No.11851303).